Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
5x + 5y = 50
10x + 6y = 84

Answer by lwsshak3(11628)

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
1) 5x + 5y = 50
2) 10x + 6y = 84
..
multiply eq 1 by -2
3) -10x-10y=-100
2) 10x + 6y = 84
add
-4y=-16
y=4
..
5x=50-5y=50-20=30
x=30/5=6
one solution, lines intersect at (6,4)